4 On January 1, 2018, Ithaca Corp. purchases Cortland Inc. bonds that have a face value of $220,000. The Cortland bonds have a stated interest rate of 5%. Interest is paid semiannually on June 30 and December 31, and the bonds mature in 10 years. For bonds of similar risk and maturity, the market yield on particular dates is as follows: (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) 5 points January 1, 2018 June 30, 2018 December 31, 2018 7.0% 8.0% 9.0% eBook Print Required 1. Calculate the price Ithaca would have paid for the Cortland bonds on January 1, 2018 (ignoring brokerage fees) 2. Prepare all appropriate journal entries related to the bond investment during 2018, assuming lthaca accounts for the bonds asa held-to-maturity investment. Ithaca calculates interest revenue at the effective interest rate as of the date it purchased the bonds. 3. Prepare all appropriate journal entries related to the bond investment during 2018, assuming that Ithaca chose the fair value option when the bonds were purchased, and that Ithaca determines fair value of the bonds semiannually. Ithaca calculates interest revenue at the effective interest rate as of the date it purchased the bonds. Complete this question by entering your answers in the tabs below Required 1 Required 2 Required 3 Prepare all appropriate journal entries related to the bond investment during 2018, assuming Ithaca accounts for the bonds as a held-to- maturity investment. Ithaca calculates interest revenue at the effective interest rate as of the date it purchased the bonds. (If no entry is required for a transaction/event, select No journal entry required in the first account field. Do not round your intermediate calculations and round your final answers to nearest whole number.) Show lessAView transaction list 1 Record the investment in bonds with a face value of $220,000, a stated interest rate of 5% and a market yield of 7%. The bonds pay interest semi-annually. ed semi 2 Record the interest revenue. 3 Record the fair value adjustment when the market yield / is 896 4 Record the interest revenue. 5 Record the fair value adjustment when the market yield Credit is 996 Note -journal entry has been entered Record entry Clear entry View general journal
Answer #1
A B C D E F G H I
2
3 Jan 1 2018
4 Face Value of the Bond $220,000
5 Coupon (Semi-Annual) 5%
6 Period (years) 10
7 Market Yield 7%
8
9 Semi-Annual coupon payment $5,500 =D4*D5/2
10 Semi-annual period 20 =D6*2
11 Semi-annual market Yield 3.50% =D7/2
12
13 Current Price of the bond will be the present value of the future cash flows.
14 PV function of excel can be used to calculate the present value of coupon and face value.
15
16 PV of bond $188,732.71 =PV(D11,D10,-D9,0)+D4/((1+D11)^D10)
17
18 Discount on Bond =Face Value -Price of the Bond
19 $31,267.29 =D4-D16
20
21 1)
22 A held to maturity investment is made by a company which it intends to hold till maturity of the investment.
23 A held to maturity investment is reported on balance sheet at its amortized cost.
24
25 Journal entry for purchase of bonds:
26 Date Account Debit Credit
27 1/1/2018 Investment in Bonds $220,000
28 Discount on Bonds $31,267
29 Cash $188,733
30
31 Year Cash Received Interest Revenue(@3.50% of BV) Amortization of Bond Discount Carrying value of Bond
32 1/1/2018 $188,733
33 6/30/2018 $5,500 $             6,605.64 $         1,105.64 $ 189,838.36
34 12/31/2018 $5,500 $             6,644.34 $         1,144.34 $ 190,982.70
35
36 2)
37 Journal entry for interest income received:
38 Date Account Debit Credit
39 6/30/2018 Cash $5,500.00
40 Discount on Bond $1,105.64
41 Interest Revenue $6,605.64
42
43 3)
44 Since the security is held to maturity, therefore it will be held till the maturity,
45 therefore no fair value adjustment is required for held to maturity security.
46
47 4)
48 Journal entry for interest income received:
49 Date Account Debit Credit
50 6/30/2018 Cash $5,500.00
51 Discount on Bond $1,144.34
52 Interest Revenue $         6,644.34
53